OpenStudy (anonymous):

Create three fractions by replacing the word 'blank' in each fraction with a variable or number. You can use the same value or three different replacement values. Explain all the work in simplifying your new addition problem. 3 over blank plus blank over a squared minus blank over a cubed

OpenStudy (anonymous):

\[\frac{ 3 }{ blank } + \frac{ blank }{ a^2 } + \frac{ blank }{ a^3 }\]

OpenStudy (anonymous):

for the first blank, i would choose 'a' for the second blank, i would choose 'a' for the third blank, i would choose 'a^2' the will create\[\frac{ 3 }{ a }+\frac{ a }{ a^{2} }+ \frac{ a^{2} }{ a^{3} }\]

OpenStudy (anonymous):

this will reduce to \[\frac{ 3 }{ a }+\frac{ 1 }{ a }+\frac{ 1 }{ a }\]

OpenStudy (anonymous):

which then reduces to \[\frac{ 5 }{ a }\]